PART ONE:

PART TWO:

PART THREE:

PART FOUR:

Gaseous methane (CH4) will react with gaseous oxygen (O2) to produce gaseous carbon dioxide (CO2 and gaseous water (H20). Suppose 13. g of methane is mixed with 13.8 g of oxygen. Calculate the minimum mass of methane that could be left over by the chemical reaction. Round your answer to 2 significant digits
A 1.32-g sample of a compound is burned in a bomb calorimeter, producing a temperature change from 20.85 °C to 27.65 °C. The heat capacity of the calorimeter is determined to be 4.84 kJ/°C. What is E (aka ДU, in kJ/g) for the combustion of this compound? Enter your answer as an integer.
Fe2O3(s) + 2 Al(s) Al2O3(s) + 2 Fe(s) ДНо 851.5 kJ/mol If 75 g Fe203 is combined with 23 g Al, what is the value of dq (kJ)? Enter your answer as an integer.
Suppose 160 mL of 8.0 M NaOH(aq) is added to 117 mL of 6.0 M HNO3(aą) to make a solution with a density of 1.1 g/mL and a specific heat of 4.1 J/gxK. If the initial temperature of both solutions is 295 K, what will be the final temperature (K)? Enter your answer as an integer. H+(aq) + OH-(aq) H2O(1) ΔΗΡ-_55.8 kJ/mol
Answer #1

Part 1

CH4(g) + 2O2(g)   \rightarrow CO2(g) + 2H2O(g)

1 mole 2 mole 1 mole 2 mole

Molecular mass of methane = 12 + (4 x 1) = 16 gm/mol

Molecular mass of oxygen, O2 = 2 x 16 = 32 gm /mol

A per the balance chemical equation,

2 mole of oxygen reacts with 1 mole of methane

or, (2 x 32) gm of oxygen reacts with 16 gm of methane

or, 13.8 gm of oxygen reacts with [16 / (2 x 32)] x 13.8 gm of methane

= 3.45 gm of methane

Available methane = 13.0 gm

Thus, methane that remains unreacted = 13.0 - 3.45 = 9.55 gm \approx 9.6 gm